Time Value of Money Problems:

Can someone please create and solve at least three time value of money problems in paragraph form. Two of the three time value of money problems should be general present value, future value, or payment problems given with an answer that includes the correct keystrokes on the calculator, as well as the correct numerical answer. The third problem must be an Amortization problem, with the correct calculator keystrokes given.
Answer #1

1- Suppose You Invest 5000 for a Year at 12% per year find the future value of the Investment

Future value = PV*(1+r)^n = 5000*(1.12)^1 = 5600

The value of investment would be 5600 after one year

2- Suppose You will receive 5000 after 1 Year, interest rate is 12% per year find the present value of the Investment

PV = future value/(1+r)^n = 5000/(1.12)^1 = 4464.28

3-

Suppose You financed a car with a loan amount of 100000 which is to be paid at 12 years period at a interest rate of 12% find the amount of yearly payment to amortize the debt

Monthlt payment to amortize the loan

amount financed/ PVAF at 12% for Year 1

100000/6.1943

16143.8742

PVAF = 1-(1+r)^-n / r

1-(1+.12)^-12/ 12%

0.7433/.12

6.1943